OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10)

OpenStudy (anonymous):

y-x=6

OpenStudy (anonymous):

m1=m2 y=x+8 m1=1 slop=y-x/y-x 1=y-10/x-4 y-10=1(x-4) y=x-6